OpenStudy (anonymous):

(x-2/x+2)+(4/x+2)+4=0

OpenStudy (anonymous):

since they are the same denominator would it be x+2 over x+2 because if u add the 4 to the -2 then its just 2

OpenStudy (nirmalnema):

x-2+4=(-4)(x+2) x+2=-4x-8 x=-2

OpenStudy (anonymous):

thanks

OpenStudy (nirmalnema):

ur welcome